Find the inverse of y=3 to the power of x +6
1
Expert's answer
2020-04-16T17:27:42-0400

Consider the function "y=3^{x+6}"


Interchange the variable "x" and "y" to obtain,



"x=3^{y+6}"

Take log of base "e" both sides as,



"ln( x)=ln( 3^{y+6})"

Use the power property of logarithm "ln( a^{m})=mln( a)" to simplify the right hand side of the equation as,



"ln( x)=(y+6 )ln( 3)"




"ln( x)=yln( 3)+6ln( 3)"




"ln( x)-6ln( 3)=yln( 3)"




"y=\frac{ln( x)-6ln( 3)}{ln( 3)}"

The sketch of the curve and its inverse is as shown in the figure below:





Here, the curve and it's inverse is symmetric about the line "y=x" .


Therefore, the inverse of the function "y=3^{x+6}" is "y=\frac{ln( x)-6ln( 3)}{ln( 3)}"